# Nonlinear particle motion and bursty periodic energy deposition in inductively coupled plasmas

### v2

@@ Abstract
where $v_{\perp,tr}=qB(t)\delta/2m_{e}$ is the maximum perpendicular velocity for the electrons to be trapped in the effective potential well (see Eq. (9)), since only the trapped electrons in $f_{0}$ contribute to the current. The velocity $v^{\prime}_{F}=E_{y}/B(t)$ contains only the $E\times B$ drift. Note that in Eq. (8) we have neglected the diamagnetic current since in our case, it is much smaller than the $E\times B$ current [84].
@@
Our case is more complicated than the standard gyrokinetic treatment [88, 89, 90] in two aspects. First, only the electrons are magnetized, so the current from electron $E_{y}\times B$ drift dominates and cannot be canceled with that of the ions. Second, the guiding magnetic field is time-varying and the drift velocity $v_{F}$ is comparable to or larger than the thermal velocity $v_{th}$, making the terms associated with $v_{F}$ dominant. Because of this, the contribution from $f_{1}$ to the current is not negligible because there are terms proportional to $N_{e}$ and $v_{F}$. In fact, $|J_{ze,1}|\sim 0.5|J_{ze,0}|$. Equation (7) for $J_{ze,1}$ contains terms involving Bessel functions. Those represent the Finite Larmor Radius (FLR) effect. Physically, the FLR effect occurs because when a particle undergoes gyromotion and a field gradient is present, the particle tends to spend more time on the low-field side than the high-field side [91, 92, 93, 89]. As a consequence, the effective field acting on the particle will be reduced, hence reducing the current. Therefore, in fact, $J_{ze,1}$ takes the sign opposite to that of $J_{ze,0}$ and acts to reduce the current. A reduced expression based on Eq. (6) is also provided in our accompanied paper (Eq. (12) of Ref. [84]). It can be used for estimating the amplitude of the current in fluid simulation codes and experimental work. Physically, the dominant influence of the RF magnetic field on particle motion causes the relationship between the electron current and the RF electric field to be nonlinear, indeed indicating the presence of a new physical regime. With the gradual increase of coil current, the plasma density increases and the electron trajectory changes to the classical type as seen in Fig. 2 (b3), with all drifts becoming negligible and the electron current taking the form predicted in existing kinetic theory $\hat{J}_{ze,old}=\omega q^{2}/m_{e}\int v_{z}/(\omega-k_{y}v_{y})\partial f_{0}/\partial\vec{v}\cdot\hat{A}d\vec{v}$ [16, 45, 84]. When neglecting all drift velocities and taking the limit $B\to 0$, one can also prove that the third term contributing to $J_{ze,1}$ in our new theory reduces to the previously known kinetic result [84].
@@
Figure 4: Comparing PIC simulations with theory for (a): $1~{}\mathrm{MHz}$ case and (b): $10~{}\mathrm{MHz}$ case. Subplots (a1) and (b1) show the electron current $J_{ze}$ and subplots (a2) and (b2) show the energy deposition rate. The comparison is made at the locations denoted by the black stars in Fig. 1 (a4) and (b4). For our new kinetic theory, we use Eq. (6) with $k_{y}=2\pi n_{y}/L_{y}$ with $n_{y}$ in the range from $1$ to $210$ (the number of cells occupied by plasma) to compute $J_{ze,1}$. The fluid approach is from Ref. [17] and the non-local kinetic calculation is from Ref. [45].
@@
Figure 4 compares the time evolution of $J_{ze}$ and of the energy deposition between the theoretical predictions and the PIC simulations. For our presently proposed theory (black) and the Tuszewski theory (red dashed, based on a fluid model without accounting for self-consistent electrostatic field) [17], we evaluate the electric field and other relevant quantities at the location marked by the black star in Fig. 1 (a4) and (b4). We then calculate the electron current from Eq. (6) and the conductivity given in Ref. [17], respectively. To carry out a calculation based on the previous non-local kinetic theory developed for high frequency ICP (blue), we additionally employ the code provided in Ref. [45] to perform a 2D simulation of the entire system, and then extract the electric field and electron current at the same location marked in Fig. 1 (a4), (b4). Note that this existing kinetic theory assumes plasma density to be uniform across the simulation domain. It is seen that the fluid theory does not predict the current and the energy deposition rate, whereas the previous kinetic theory performs well only in the $f=10~{}\mathrm{MHz}$ case. Our new theory, on the other hand, displays a much better agreement for both the current and the energy deposition at low frequency. We identify the new regime dominated by periodic bursty energy deposition as the “periodic burst regime”. The above comparison is performed at a single location because the electrons are trapped in the strong effective potential and the plasma response is nearly local. When averaging over the entire simulation domain, the jet-like current, comprised of the released trapped electrons (see Fig. 2 (a3)), contributes around $70\%$ to the total energy deposition.
@@
Parameter Scan–Next, we perform a parameter scan over $I_{coil}$ and $\omega$ to identify the onset boundary for the periodic burst regime. Figure 5 shows the average energy deposition, average ion density and average electron temperature for different values of the driving waveform parameters. We see that at lower frequencies there is a distinct jump in energy deposition from $10^{-2}W/cm^{3}$ to $10^{0}W/cm^{3}$, indicating a transition from the periodic burst regime to the anomalous (non-local) skin effect regime. A similar jump is observed for the ion density. Physically, this transition occurs because in the periodic burst regime, the electric field and current are nearly out of phase during most of the RF period. This is because the strong RF magnetic field makes plasma response nearly local. Only in the phase of the electron jet the energy deposition is seen to become significant. Therefore, the effective energy deposition becomes much lower than in the anomalous skin effect regime, making the electron density also much lower. Therefore, based on this understanding of particle dynamics, the criterion for transition between the bursty regime and the usual anomalous non-local regime can be produced by equating twice the electron gyro-radius with the anomalous field skin depth:
$2\frac{m_{e}v_{th}}{qB}=\left(\frac{v_{th}c^{2}}{\omega\omega^{2}_{pe}}\right)^{1/3}.$ (9)
@@
When the diameter of the electron gyro-circle is larger than the skin depth, the electrons are unlikely to undergo a full gyro-cycle, and are therefore no longer being trapped in the RF effective potential. Based on Eq. (9), an estimate of the plasma density at the threshold of the periodic burst regime is obtained. The cases with different driving frequencies will have different transition values for the plasma density, as indicted in Fig. 5 (b) by the dashed lines. One expects to obtain the periodic burst regime in the region located below the dashed lines. A reasonable agreement is found, with the red and blue stars in (b) denoting the transition point, indicating the validity of Eq. (9) for finding the boundary between the periodic burst regime and the anomalous non-local regime.
@@
Figure 5: The solid lines show, in a steady state, (a) energy deposition rates, averaged over the entire simulation domain and RF period, (b) time-averaged ion density at the probe location denoted by the black star in Fig. 1 (a4), and (c) time-averaged electron temperature at the probe location denoted by the black star in Fig. 1 (a3), for different values of the coil current and driving frequency. The dashed lines denote the onset threshold for the periodic burst regime, which follows the scaling law of $N_{e}\sim I_{coil}^{3}/\omega T_{e}$ according to Eq. (9). Each dashed line denotes the boundary for same frequency as the solid line of the same color in (b). As an estimate, the relevant magnetic field and temperature data for calculating the density (denoted by dashed lines) based on Eq. (9) are evaluated at the black star locations in Fig. 1 (a4) and (b4). The red and blue stars in (b) denote the transition points between the two regimes.
@@
Conclusions–We identify a periodic burst energy deposition regime in low-frequency ICP discharges where the electrons are magnetized and their dynamics is dominated by bouncing in the effective potential of the electromagnetic and electrostatic fields through most of the RF period, so that the electron response becomes nearly local in space. When the RF magnetic field becomes weak, the electrons demagnetize, forming a jet-like current which propagates quickly into the plasma and accounts for a sizable fraction of the energy deposition rate. A new kinetic theory is proposed to estimate the electron current, which is in good agreement with particle-in-cell simulations. The condition in the driving waveform parameter space for the transition from anomalous skin effect regime (unmagnetized electrons) to the periodic burst regime is identified as a result of analyzing the data produced in a series of 2D simulations and is also analytically derived. A significant jump in the plasma density is observed during such transition, akin to what was observed in the E-H mode transition [94, 95, 20, 83, 96]. Therefore, our finding also offers a new possible explanation to the density jumps in the transition between E-mode and H-mode in low-frequency, low pressure ICP. We propose an estimate for the amplitude of the current density in the skin layer (Eq. (6)) and a criterion for transition to the periodic burst regime (Eq. (9)) in a form that can be verified in future experiments.

### v3

@@ Abstract
where $v_{\perp,tr}=|qB(t)|\delta/2m_{e}$ is the maximum perpendicular velocity for the electrons to be trapped by the magnetic field, since only the trapped electrons in $f_{0}$ contribute to the current. And $v^{\prime}_{F}=E_{y}/B(t)$. In Eq. (4) we have neglected the diamagnetic current since it is much smaller than the $E\times B$ current.
@@
Our case is more complicated than the standard gyrokinetic treatment [90, 91, 92] in two aspects. First, only the electrons are magnetized, so the current from electron $E_{y}\times B$ drift dominates and cannot be canceled with that of the ions. Second, the guiding magnetic field is time-varying and the drift velocity $v_{F}$ is comparable to or larger than the thermal velocity $v_{th}$, making the terms associated with $v_{F}$ dominant. Because of this, the contribution from $f_{1}$ to the current is not negligible because there are terms proportional to $N_{e}$ and $v_{F}$. Equation (3) for $J_{ze,1}$ contains Bessel functions. Those represent the Finite Larmor Radius (FLR) effect. Physically, the FLR effect occurs because when a particle undergoes gyro-motion and a field gradient is present, the particle tends to spend more time on the low-field side than the high-field side [93, 94, 95, 91]. Therefore effectively, the field acting on the electrons will be reduced, hence reducing the current. Therefore, $J_{ze,1}$ takes the sign opposite to that of $J_{ze,0}$ and acts to reduce the current. The dominance of the RF magnetic field on particle motion significantly changes the electron current (see simplified formula Eq. (12) in Ref. [86]), indicating the presence of a new physical regime. With the gradual increase of coil current, the plasma density increases and the electron trajectory changes to the classical type as seen in Fig. 2 (b3), with all drifts becoming negligible and the electron current taking the form in existing kinetic theory $\hat{J}_{ze,old}=\omega q^{2}/m_{e}\int v_{z}/(\omega-k_{y}v_{y})\partial f_{0}/\partial\vec{v}\cdot\hat{A}d\vec{v}$ [16, 47].
@@
Figure 3: Comparing PIC simulations with theory for (a): $1~{}\mathrm{MHz}$ case and (b): $10~{}\mathrm{MHz}$ case. Subplots (a1) and (b1) show the electron current $J_{ze}$ and subplots (a2) and (b2) show the energy deposition rate. The comparison is made at the locations denoted by black stars in Fig. 1 (a4) and (b4). For our new kinetic theory, we use Eq. (2) with $k_{y}=2\pi n_{y}/L_{y}$ with $n_{y}$ from $1$ to $210$ (number of cells occupied by plasma) to compute $J_{ze,1}$. The fluid approach is from Ref. [17] and the previous non-local kinetic calculation is from Ref. [47].
@@
Figure 3 compares the time evolution of $J_{ze}$ and energy deposition between the theoretical predictions and the PIC simulations. For our new theory (black) and the Tuszewski theory (red dashed, based on fluid model) [17], we evaluate the electric field and other relevant quantities at the location marked by the black stars in Fig. 1 (a4) and (b4). We then calculate the electron current from Eq. (2) and the conductivity given in Ref. [17], respectively. To carry out a calculation based on the previous non-local kinetic theory developed for high frequency ICP (blue), we employ the code provided in Ref. [47] to perform a 2D simulation of the entire system, and then extract the electric field and electron current at the same location marked in Fig. 1 (a4), (b4). This existing kinetic theory assumes uniform plasma density. We see that the fluid theory does not predict the current and the energy deposition rate, whereas the previous kinetic theory performs well only for $f=10~{}\mathrm{MHz}$ case. Our new theory, on the other hand, displays a much better agreement for both the current and the energy deposition at low frequency. We identify the new regime dominated by periodic bursty energy deposition as the “periodic burst regime”. When averaging over the simulation domain, the jet-like current contributes around $70\%$ to the total energy deposition.
@@
Parameter Scan–We perform a parameter scan over $I_{coil}$ and $\omega$ to identify the onset boundary for the periodic burst regime. Figure 4 shows the average energy deposition, ion density, and electron temperature for different values of the driving waveform parameters. We see that at lower frequencies there is a distinct jump in energy deposition, indicating a transition from the periodic burst regime to the anomalous (non-local) skin effect regime. A similar jump is observed for the ion density. This transition occurs because in the periodic burst regime, the electric field and current are nearly out of phase during most of the RF period, since the strong RF magnetic field makes plasma response nearly local. Only in the phase of electron jet will the energy deposition become significant. Therefore, the energy deposition becomes lower than in the anomalous skin effect regime, making the electron density also much lower. Based on this understanding of particle dynamics, the criterion for transition between the bursty regime and the anomalous non-local regime is estimated by equating twice the electron gyro-radius with the anomalous field skin depth [86]
$2\frac{m_{e}v_{th}}{|qB|}=\left(\frac{v_{th}c^{2}}{\omega\omega^{2}_{pe}}\right)^{1/3}.$ (5)
@@
When the diameter of the electron gyro-circle is larger than the skin depth, the electrons are unlikely to undergo a full gyro-cycle, and are therefore no longer being confined by the magnetic field. Based on Eq. (5), an estimate of the plasma density at the threshold of the periodic burst regime is obtained. The cases with different driving frequencies will have different transition values for the plasma density, as indicated in Fig. 4 (b) by the dashed lines. One expects to obtain the periodic burst regime in the region located below the dashed lines. A reasonable agreement is found, indicating the validity of Eq. (5) for finding the boundary between the two regimes.
@@
Figure 4: The solid lines show, in a steady state, (a) energy deposition rates, averaged over the entire simulation domain and RF period, (b) time-averaged ion density at the probe location denoted by the black star in Fig. 1 (a4), and (c) time-averaged electron temperature at the probe location denoted by the black star in Fig. 1 (a3), for different values of the coil current and driving frequency. The dashed lines denote the onset threshold for the periodic burst regime, which follows the scaling law of $N_{e}\sim I_{coil}^{3}/\omega T_{e}$ according to Eq. (5). Each dashed line denotes the boundary for the solid line of the same color in (b). The relevant magnetic field and temperature data for calculating the dashed lines is evaluated at the black stars in Fig. 1 (a4) and (b4). The red and blue stars in (b) denote the transition points between the two regimes.
@@
Conclusions–We identify a periodic burst energy deposition regime in low-frequency ICP discharges where the electrons are magnetized, so that the electron response becomes nearly local during most of the RF period. When the RF magnetic field becomes weak, the electrons demagnetize, forming a jet-like current propagating quickly into the plasma and accounting for a sizable fraction of the energy deposition. A new kinetic theory is proposed to estimate the electron current, showing good agreement with PIC simulations. The condition for the transition from anomalous skin effect regime (unmagnetized electrons) to the periodic burst regime is both identified in simulations and given analytically. A significant jump in the plasma density is observed during such transition, akin to what was observed in the E-H mode transition [96, 97, 20, 85, 98]. Our finding also offers a new possible explanation for the density jumps in the transition between E-mode and H-mode in low-frequency, low pressure ICP. We propose an estimate for the current density in the skin layer (Eq. (2)) and a criterion for transition to the periodic burst regime (Eq. (5)) that can be verified in future experiments.